a solution contains 3% salt. How much water should be added to 20oz. of this solution to make a 1.2% solution?
Answer by stanbon(75887)

a solution contains 3% salt. How much water should be added to 20oz. of this solution to make a 1.2% solution?
-----
Equation:
salt + salt = salt
0.03*20 + 0*x = 0.012(20+x)
0.6 + 0 = 0.24 + 0.012x
-------
0.012x = 0.36
-------
x = 0.360/0.012
x = 30 oz (amt of water needed)
--------------
Cheers,
Stan H.